OpenStudy (anonymous):

y=x^3-5x^2-x+5 determine excluded regions.

OpenStudy (anonymous):

I really just need help grouping.

OpenStudy (anonymous):

i only get confused by the constant term .. else I would just pull out x

OpenStudy (anonymous):

\[y=x ^{3}-5x ^{2}-x+5\]

OpenStudy (anonymous):

Are u trying to factor it

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

take x-1 common

OpenStudy (anonymous):

y=x^3-5x^2-x+5 =x^3-x^2-4x^2+4x-5x+5 =x^2(x-1)-4x(x-1)-5(x-1)

OpenStudy (anonymous):

Now, can u do it from here.

OpenStudy (anonymous):

\[y=x ^{2}(x-5)-x+5\]

OpenStudy (anonymous):

Yes that also works

OpenStudy (anonymous):

hmm, how would i group the 2nd term -x+5? or do i not need to do anything there?

OpenStudy (anonymous):

y=x^2(x-5)-1(x-5)

OpenStudy (anonymous):

\[y=x ^{2}(x-5)+(x-5)\]?

OpenStudy (anonymous):

ahh so not pull out anything.. just group x-5

OpenStudy (anonymous):

\[(x ^{2}-1)(x-5)= (x-1)(x+1)(x-5)\]?

OpenStudy (anonymous):

Yes very correct

OpenStudy (anonymous):

woo awesome. thanks for your help, very much appreciated.

OpenStudy (anonymous):

Welcome